# VISO-Grasp: Vision-Language Informed Spatial Object-centric 6-DoF Active View Planning and Grasping in Clutter and Invisibility

### v1

@@ Abstract
We propose VISO-Grasp, a novel vision-language-informed system designed to systematically address visibility constraints for grasping in severely occluded environments. By leveraging Foundation Models (FMs) for spatial reasoning and active view planning, our framework constructs and updates an instance-centric representation of spatial relationships, enhancing grasp success under challenging occlusions. Furthermore, this representation facilitates active Next-Best-View (NBV) planning and optimizes sequential grasping strategies when direct grasping is infeasible. Additionally, we introduce a multi-view uncertainty-driven grasp fusion mechanism that refines grasp confidence and directional uncertainty in real-time, ensuring robust and stable grasp execution. Extensive real-world experiments demonstrate that VISO-Grasp achieves a success rate of $87.5\%$ in target-oriented grasping with the fewest grasp attempts outperforming baselines. To the best of our knowledge, VISO-Grasp is the first unified framework integrating FMs into target-aware active view planning and 6-DoF grasping in environments with severe occlusions and entire invisibility constraints.
@@ I Introduction
Robotic grasping in unstructured, cluttered environments remains a significant challenge, particularly for executing target-oriented grasps under partial or complete occlusions [1]. Humans instinctively overcome occlusion during targeted searches by adjusting their viewpoints and intuitive reasoning about spatial relationships to infer potential target locations. In comparison, leveraging the recent success of deep neural networks and contributions to large-scale training data [2, 3, 4], existing learning-based robotic grasp detectors [1] rely primarily on static viewpoint observations, assuming sufficient visibility of the target object from single or pre-defined multi-view perspectives[5, 6, 7]. This constraint reduces flexibility in scenarios where the target object is highly occluded or entirely unobservable.
@@
To this end, we propose VISO-Grasp, a Vision-language Informed Spatial Object-centric grasping framework that integrates off-the-shelf Foundation Models (FMs) [11, 12, 13] with active vision and occlusion-aware $\mathrm{SE(3)}$ grasp planning, which leverages the inherent prior of VLMs to achieve human-like decision-making. In contrast to recent works [14, 15, 16], which focus on planar 4-DoF grasp reasoning with VLMs while overlooking their fundamental limitations in accurate 3D spatial understanding and real-time execution, VISO-Grasp introduces evolving spatial reasoning that continuously integrates spatial and occlusion relationships. By incorporating an online grasp fusion mechanism, our approach dynamically refines target visibility and substantially improves grasp efficiency in heavy occlusions. Specifically, when direct grasping is infeasible, a sequential grasp strategy guided by VLMs is employed, leveraging the observable relational hierarchy to determine the optimal grasp sequence for decluttering and removing potential obstructions. In summary, our main contributions are threefold:
@@
- Uncertainty-driven online grasp fusion, a real-time grasp fusion framework that integrates multi-view grasp predictions through pose-centric categorization and real-time Bayesian updates. This leverages a pre-trained uncertainty-aware grasp generator [17], trained exclusively on simulated data, to enhance grasp robustness and reliability.
@@ II-B FMs in Robotic Perception and Planning
Foundation Models (FMs), including VLMs, have significantly improved robotic perception and planning by zero-shot object detection [27, 28], multi-modal reasoning [29, 16, 30], and decision-making [31, 32], reducing reliance on task-specific data collection while enhancing adaptability in real-world environments [33, 34]. Recent works leverage FMs to enable open-vocabulary perception, integrating zero-shot segmentation for novel object recognition  [15, 14], while uncertainty-aware frameworks refine scene understanding by dynamically calibrating perception and decision uncertainties [35]. In robotic planning, multimodal models facilitate goal-conditioned decision-making by aligning spatial and semantic representations [36], while vision-language systems enhance strategic scene exploration and adaptive reasoning in unstructured scenarios [37]. Yet, most FM-driven approaches still rely on 2D representations, lacking explicit 3D spatial priors and depth reasoning [38]. Furthermore, they assume minimal occlusions and primarily operate under SE(2) planar constraints, with limited real-time viewpoint adaptation and active perception to dynamically refine scene understanding. To overcome these limitations, we propose the first FMs-driven framework that unifies multi-view 3D spatial reasoning and active viewpoint adaptation for occlusion-aware 6-DoF grasping, enabling dynamic scene refinement and grasp execution in cluttered environments.
@@ III Problem Formulation
To achieve target-oriented grasping in these scenarios, we incorporate FMs [11, 13, 12] for semantic scene understanding and adaptive view planning, allowing zero-shot generalization without task-specific training. In extreme cases such as complete occlusions or no valid grasp poses available on the target, the VLM either determines an optimal grasp execution sequence from current observed objects $\mathcal{O}^{\prime}$ to prioritize the removal of occluders, or infers the potential occluders by the VLM with VICL from historical observations.
@@ IV-D2 Contact-grasp Fusion
Here $\kappa^{t}_{\mathbf{c}}$ represents the precision on the observed mean likelihood $\mu_{\mathbf{c}}$. We refer interesting readers for the theoretical background to [17].

### v2

@@ Abstract
We propose VISO-Grasp, a novel vision-language-informed system designed to systematically address visibility constraints for grasping in severely occluded environments. By leveraging Foundation Models (FMs) for spatial reasoning and active view planning, our framework constructs and updates an instance-centric representation of spatial relationships, enhancing grasp success under challenging occlusions. Furthermore, this representation facilitates active Next-Best-View (NBV) planning and optimizes sequential grasping strategies when direct grasping is infeasible. Additionally, we introduce a multi-view uncertainty-driven grasp fusion mechanism that refines grasp confidence and directional uncertainty in real-time, ensuring robust and stable grasp execution. Extensive real-world experiments demonstrate that VISO-Grasp achieves a success rate of $87.5\%$ in target-oriented grasping with the fewest grasp attempts outperforming baselines. To the best of our knowledge, VISO-Grasp is the first unified framework integrating FMs into target-aware active view planning and 6-DoF grasping in environments with severe occlusions and entire invisibility constraints. Code is available at: https://github.com/YitianShi/vMF-Contact
@@ I Introduction
Robotic grasping in unstructured, cluttered environments remains a significant challenge, particularly for executing target-oriented grasps under partial or complete occlusions [1]. Humans instinctively overcome occlusion during targeted searches by adjusting their viewpoints and intuitive reasoning about spatial relationships to infer potential target locations. In comparison, leveraging the recent success of deep neural networks and contributions to large-scale training data [2, 3, 4, 5], existing learning-based robotic grasp detectors [1] rely primarily on static viewpoint observations, assuming sufficient visibility of the target object from single or pre-defined multi-view perspectives[6, 7, 8]. This constraint reduces flexibility in scenarios where the target object is highly occluded or entirely unobservable.
@@
To this end, we propose VISO-Grasp, a Vision-language Informed Spatial Object-centric grasping framework that integrates off-the-shelf Foundation Models (FMs) [12, 13, 14] with active vision and occlusion-aware $\mathrm{SE(3)}$ grasp planning, which leverages the inherent prior of VLMs to achieve human-like decision-making. In contrast to recent works [15, 16, 17, 18], which focus on planar 4-DoF grasp reasoning with VLMs while overlooking their fundamental limitations in accurate 3D spatial understanding and real-time execution, VISO-Grasp introduces evolving spatial reasoning that continuously integrates spatial and occlusion relationships. By incorporating an online grasp fusion mechanism, our approach dynamically refines target visibility and substantially improves grasp efficiency in heavy occlusions. Specifically, when direct grasping is infeasible, a sequential grasp strategy guided by VLMs is employed, leveraging the observable relational hierarchy to determine the optimal grasp sequence for decluttering and removing potential obstructions. In summary, our main contributions are threefold:
@@
- Uncertainty-driven online grasp fusion, a real-time grasp fusion framework that integrates multi-view grasp predictions through pose-centric categorization and real-time Bayesian updates. This leverages a pre-trained uncertainty-aware grasp generator [19], exclusively on simulated data, to enhance grasp robustness and reliability.
@@ II-B FMs in Robotic Perception and Planning
Foundation Models (FMs), including VLMs, have significantly improved robotic perception and planning by zero-shot object detection [29], multi-modal reasoning [30, 17, 31], and decision-making [32, 33], reducing reliance on task-specific data collection while enhancing adaptability in real-world environments [34, 35]. Recent works leverage FMs to enable open-vocabulary perception, integrating zero-shot segmentation for novel object recognition  [16, 15], while uncertainty-aware frameworks refine scene understanding by dynamically calibrating perception and decision uncertainties [36]. In robotic planning, multimodal models facilitate goal-conditioned decision-making by aligning spatial and semantic representations [37], while vision-language systems enhance strategic scene exploration and adaptive reasoning in unstructured scenarios [38]. Yet, most FM-driven approaches still rely on 2D representations, lacking explicit 3D spatial priors and depth reasoning [39]. Furthermore, they assume minimal occlusions and primarily operate under SE(2) planar constraints, with limited real-time viewpoint adaptation and active perception to dynamically refine scene understanding. To overcome these limitations, we propose the first FMs-driven framework that unifies multi-view 3D spatial reasoning and active viewpoint adaptation for occlusion-aware 6-DoF grasping, enabling dynamic scene refinement and grasp execution in cluttered environments.
@@ III Problem Formulation
To achieve target-oriented grasping in these scenarios, we incorporate FMs [12, 14, 13] for semantic scene understanding and adaptive view planning, allowing zero-shot generalization without task-specific training. In extreme cases, such as complete occlusions or no valid grasp poses available on the target, the VLM either determines an optimal grasp execution sequence from current observed objects $\mathcal{O}^{\prime}$ to prioritize the removal of occluders, or infers the potential occluders by the VLM with VICL from historical observations.
@@ IV-D2 Contact-grasp Fusion
Here $\kappa^{t}_{\mathbf{c}}$ represents the precision on the observed mean likelihood $\mu_{\mathbf{c}}$. We refer interested readers for the theoretical background to [19].